# Qumode Tensor Networks for False Vacuum Decay in Quantum Field Theory

### v1

@@ II.6 Initial state preparation
$\Delta{q}_{n}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{\alpha=0}^{N-1}\frac{1}{\sqrt{2N}}\sqrt{\left(\frac{1}{\omega}-\frac{1}{\omega_{\alpha}}\right)}\leavevmode\nobreak\ \cos(2\pi n\alpha/N)\leavevmode\nobreak\ {\eta}_{n,\alpha}\leavevmode\nobreak\ ,$
$\Delta p_{n}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{\alpha=0}^{N-1}\sqrt{\frac{\omega_{\alpha}-\omega}{2N}}\,\cos(2\pi n\alpha/N)\leavevmode\nobreak\ {\eta}^{\prime}_{n,\alpha}\leavevmode\nobreak\ ,$ (49)
where $\eta_{n}$ and $\eta_{n}^{\prime}$ are standard normal vectors, and the angular frequencies $\omega$ and $\omega_{\alpha}$ are given in Eqs. (3) and (71) respectively. (The shift $\Delta p$ is of course implemented by acting on the wave function of each qumode with the operator $e^{i\Delta p_{n}\,\hat{q}_{n}}$.)
III False vacuum decay in (1+1)-dimensional QFT
III.1 The objective: tunnelling solutions
@@
$\partial^{2}_{r}\varphi+r^{-1}\partial_{r}\varphi\leavevmode\nobreak\ =\leavevmode\nobreak\ \partial_{\phi}{\mathscr{V}}\leavevmode\nobreak\ .$ (50)
This now resembles the classical equation of motion of a ball rolling in the inverted potential $-\mathscr{V}$, with $r$ playing the role of time and with the $r^{-1}\partial_{r}\varphi$ term providing a friction that decays with time. Thus, the solution with the correct boundary conditions can be found by starting the ball close to the true vacuum at $\varphi\approx\varphi_{0}$ at ‘time’ $r=0$ (corresponding to the centre of the instanton) and letting it come to a halt at $r\to\infty$ precisely at the false vacuum at $\varphi=-\varphi_{0}$. This requires a fine-tuning of the starting value of $\varphi$, numerically solving the one dimensional differential equation in Eq. (50), and employing the overshoot-undershoot method described in Ref. [2] to satisfy the boundary conditions. The result for the quartic potential in Fig. 2 with $\lambda=0.5$, $\varphi_{0}=2$ and $\varepsilon=0.02$ is the bubble profile shown in Fig. 11. Although it is not possible to find a closed form for the radially symmetric solution, it can be very accurately modelled by the following function:
$\varphi(r)\leavevmode\nobreak\ \approx\leavevmode\nobreak\ -\varphi_{0}\,\tanh\beta(r-r_{0})$ (51)
@@
$\varphi(x,t)\leavevmode\nobreak\ \approx\leavevmode\nobreak\ -\,\varphi_{0}\,\tanh\beta(\sqrt{x^{2}-t^{2}}-r_{0})\leavevmode\nobreak\ .$ (52)
@@
$\varphi(x,t>|x|)\leavevmode\nobreak\ \approx-\varphi_{0}\,(1+2e^{-2\beta r_{0}}\cos{2\beta|t^{2}-x^{2}|})\leavevmode\nobreak\ .$ (53)
@@
[…]
@@
$r_{c}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{S_{1}}{2\varphi_{0}\varepsilon}\leavevmode\nobreak\ .$ (55)
@@
As mentioned in the introduction the rate of bubble nucleation per unit space-time volume is exponentially sensitive to the Euclidean action associated with the tunnelling solution. Specifically, in the thin-wall approximation it follows the Arrhenius-like behaviour $\Gamma\leavevmode\nobreak\ =\leavevmode\nobreak\ \kappa\,e^{-S_{1}}$, where $\kappa$ is a prefactor that accounts for quantum fluctuations around the instanton solution (which is notoriously difficult to evaluate). This exponential suppression means that even small increases in $S_{1}$ can lead to a dramatic decrease in the nucleation rate. Thus the value of $S_{1}$ effectively encodes how “difficult” it is for the system to tunnel from the false vacuum to the true vacuum, with lower $S_{1}$ corresponding to exponentially higher nucleation probabilities and faster phase transitions.
This then is the paradigm for false vacuum decay that was laid down in Refs. [2, 3]. Before considering the equivalent evolution on our qumode lattice, we should emphasize that there are many subtle issues surrounding the dynamics of these processes which have been discussed subsequently, and which are not fully understood: for example, whether or not the bubble “runs away” to light speed or instead reaches a terminal velocity depends on the friction generated by its interactions with the medium it is travelling through, and is a complicated and model-dependent question (see [75, 76, 77, 78] and for a recent review see [79]). This of course is one of the motivations for studying the process numerically.
@@
III.2 Results: tunnelling in (1+1)-dimensional scalar theory
To validate the model, we simulate the real-time dynamics of the system in two distinct scenarios. First, we consider “pre-formed” bubbles, initialising the system with a true-vacuum bubble with an initial radius $r_{0}$. Second, we study phase transitions in the full system by examining bubble nucleation arising from quantum fluctuations.
III.2.1 Isolated bubbles: expansion and collapse
@@
Figure 12: In 13(a) we show a collapsing subcritical bubble in the Pöschl-Teller potential of Eq. (26) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\alpha=1$, $\gamma=3/2$ and $\varepsilon=0.25$. Fig. 13(b) shows an expanding bubble in the potential of Eq. (24) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\lambda=1$, and $\varepsilon=0.1$, which shows the bubble walls following a hyperbolic trajectory that reaches a terminal speed of 0.72. Note that we do not impose periodic boundary conditions, and the bubble walls “bounce” off the sides of the lattice elastically.
@@
To begin, we examine the real-time dynamics of pre-formed bubbles with an initial radius, $r_{0}$. Figure 12 presents the results from the qumode tensor network for bubbles of different initial radii. The first scenario, shown in Fig. 13(a), considers the collapse of a subcritical bubble in the Pöschl-Teller potential from Eq. (26) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\alpha=1$, $\gamma=3/2$ and $\varepsilon=0.25$. In this example, we observe that the bubble fails to reach the critical radius. There are two interesting aspects of the collapse. The first phenomenon that can be seen has to do with the way the collapse occurs. We can clearly see the emission of several wavepackets from the walls. Averaged over time these emissions would generate friction on the walls. Here, however, we see that such emission (together with momentum conservation) is a necessary component of bubble collapse. This is because in one space-dimension there is no surface tension acting to collapse the bubble. In other words in one space-dimension the energy contained within the bubble walls is independent of the size of the bubble. The second interesting aspect of Fig. 13(a) is related. It is the fact that the final moments of collapse, in which the whole bubble “flops” back to the false vacuum, would correspond to super-luminal wall velocity. This in turn implies that the entanglement on the qumode lattice must spread over distances of at least the radius of the collapsing bubble at that instant in time, in order for the collapse to happen.
@@
Figure 13(b) shows the expansion of a bubble in the potential from Eq. (24) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\lambda=1$ and $\varepsilon=0.1$. In contrast with Fig. 13(a), this bubble is supercritical and we observe the phenomena outlined in Sec. III.1. As the bubble’s radius is greater than the critical bubble radius, the walls expand. The bubble expansion is entirely driven by internal pressure caused by the energy gain from increasing the volume of the true vacuum. This acceleration is relatively local, and long-range entanglement plays very little roll in the behaviour. As a result, the walls accelerate along hyperbolic trajectories which asymptote to a constant terminal velocity. However the terminal velocity is subluminal ($v_{\rm term}\approx 0.72$), due to the friction mentioned above.
@@
Note that we do not impose periodic boundary conditions and this results in the bubble wall colliding elastically with the sides of the lattice. We can in principle use this phenomenon to measure the mass of the bubble walls. We observe that after the bounce the walls are projected a distance of $d_{\rm bounce}\approx 20$ into the bulk of the lattice before being repelled back by the pressure, $\varepsilon$. (This relatively short distance also tells us that friction has indeed resulted in a terminal subluminal speed.) We may make an rough estimate for the mass of the wall by neglecting the effect of friction after the bounce and taking a non-relativistic approximation. Equating the initial kinetic energy with the maximum potential energy due to the pressure gives
[…]
III.2.2 Phase transitions

### v2

@@ III.1 Approximating the ground state by sampling
$\Delta{q}_{n}\penalty 10000\$$=\penalty 10000\ \sum_{\alpha=0}^{N-1}\frac{1}{\sqrt{2N}}\sqrt{\left(\frac{1}{\omega}-\frac{1}{\omega_{\alpha}}\right)}\penalty 10000\ \cos(2\pi n\alpha/N)\penalty 10000\ {\eta}_{n,\alpha}\penalty 10000\ ,$
$\Delta p_{n}\penalty 10000\$$=\penalty 10000\ \sum_{\alpha=0}^{N-1}\sqrt{\frac{\omega_{\alpha}-\omega}{2N}}\,\cos(2\pi n\alpha/N)\penalty 10000\ {\eta}^{\prime}_{n,\alpha}\penalty 10000\ .$ (63)
The upshot is that we can mimick the required long-range correlations by setting the qumodes in their “false vacuum ground states” and then augmenting them by random sampling the classical distributions of position and momentum fluctuations presented in Eq. (63), where the angular frequencies $\omega$ and $\omega_{\alpha}$ are found from Eqs. (3) and (95) respectively. (The shift $\Delta p$ is of course implemented by acting on the wave function of each qumode with the operator $e^{i\Delta p_{n}\,\hat{q}_{n}}$.)
III.2 Imaginary-time TEBD determination of the initial state
We now discuss the second option for determination of the QFT vacuum. As mentioned this option requires no compromises and is a genuine QFT initial state preparation for vacuum tunnelling. The method works by Hamiltonian evolution of the QFT through imaginary time using TEBD following the techniques discussed in Refs. [77, 78, 79]. That is we begin with the system as a product of SHO ground states, in the lifted false vacuum of Figure 4. We then perform the same Trotterized Hamiltonian evolution that we would perform for the simulation itself, but through imaginary time. The MPS will of course immediately begin to attenuate because such an evolution is non-unitary. Therefore after each Trotter step it is renormalized by scaling. The crucial point that makes the trick work is that the rate of attenuation is proportional to the energy, and over many Trotter steps the lowest energy eigenstate dominates. Thus the system is expected to converge quite rapidly to the desired QFT false vacuum state.
The great advantage of the method is precisely that it is not physical. There is no concern about for example maintaining adiabaticity, and we do not need to worry if the system gets artificially “kicked” by relatively large Trotter steps, because ultimately all these perturbations will be attenuated away.
In order to test whether the mechanism has worked we can make measurements of the same covariance matrix that we discussed in the previous section. That is we determine
$C(k_{\alpha})\penalty 10000\ =\penalty 10000\ \sum_{n,m}C_{nm}U_{\alpha n}U^{\dagger}_{m\alpha}\penalty 10000\ ,$ (64)
by measuring the matrix
$C_{nm}\penalty 10000\ =\penalty 10000\ \langle(\hat{q}_{n}-\langle q_{n}\rangle)(\hat{q}_{m}-\langle q_{m}\rangle)\rangle$ (65)
on the lattice, where $\langle q_{n}\rangle\approx\varphi_{\rm fv}$.
Ideally one would expect to find
$C(k_{\alpha})\penalty 10000\ =\penalty 10000\ \frac{1}{2\omega_{\alpha}}\penalty 10000\ ,$ (66)
as we have already discussed.
However, a fit using Eq. (95) with the exact mass deduced from the potential in $\omega_{\alpha}$ is not good. It turns out that for an accurate fit we need two parameters in $C(k_{\alpha})$. One of these parameters is an effective mass which we shall call $m$ which is inserted into the dispersion relation Eq. (95) in place of the ideal mass deduced from the potential which would have $m=\omega$. The second parameter is a constant $C_{0}$ which corresponds to an effective extra $k_{\alpha}$ independent on-site variance. Thus we fit
$C(k_{\alpha})\penalty 10000\ =\penalty 10000\ \frac{1}{2\omega_{\alpha}(m)}\penalty 10000\ +\penalty 10000\ C_{0}\penalty 10000\ .$ (67)
We will investigate the meaning of the two parameters $m$ and $C_{0}$ shortly. For the moment let us present the fits, which are shown in Figs. 11 and 12 for the unit mass SHO potential with
${\mathscr{V}}(\varphi)\penalty 10000\ =\penalty 10000\ \frac{1}{2}(\varphi-\varphi_{\rm fv})^{2}\penalty 10000\ ,$ (68)
and for the PT potential of Eq. (26) respectively, lifted and with parameters ($\alpha=1,\gamma=\frac{\sqrt{5}-1}{2}$) chosen such that it also has $\omega=1$. In this case we are taking 20 quadrature points for our qumodes (i.e. they are approximated by a local Hilbert space of dimension 20). The best fit is $C_{0}\penalty 10000\ =\penalty 10000\ -0.016$ for both theories, and in fact regardless of the mass $m$. Meanwhile the best fit mass is $m=1.0$ for the SHO potential in Fig. 11, and $m=0.72$ for the PT example shown in Fig. 12. Clearly the imaginary-time TEBD method has reproduced the desired QFT vacuum entirely as expected, modulo these two parameters.
What, then, is the meaning of these two parameters, and why are they not equal to the ideal case, which would be $m=\omega$ and $C_{0}=0$? The answer is to do with our discretization of the local Hilbert space, i.e. the discrete and finite domain of $q_{n}$ quadratures in our qumodes. We can demonstrate this in several ways.
First let us consider the mass $m$. It is arguably more consistent to use an effective mass-squared derived from the double derivative of a fit to the discretized potential ${\mathscr{V}}(q_{n})$, rather than one derived from the double derivative of the potential itself. After all, this is more indicative of the mass-squared that is “seen” by the QFT in our simulation. For the SHO potential this gives a value of $m\approx 1$, while for the PT potential this gives a mass $m\approx 0.77$, both very close to the values that we had to insert into the dispersion relation in order to fit the observed spectrum $\omega(k_{\alpha})$. This is very encouraging, and indeed this effective mass continues to track the curvature in the fit to the discretized potential rather than that of the ideal potential when we consider other models and when we vary the parameters.
The parameter $C_{0}$ has a less immediately obvious interpretation. From Eq. (64) we see that it corresponds to a $\delta$-term in $C_{nm}$ of the form $C_{nm}\supset C_{0}\delta_{nm}$. Such a term could conceivably be due to the local potential in an interacting theory, to the lattice spacing (i.e. finite $a$) or again be due to the finite local Hilbert space. (Note that it could not arise due to lack of convergence as that would be $k$ dependent.)
Figure 11: Comparison of the measured covariance matrix of the QFT vacuum versus the theoretical expressions with fit parameters $m=1$ and $C_{0}=-0.016$, for the SHO potential centred around the origin with $\omega=1$.
Figure 12: Comparison of the covariance matrix of the QFT vacuum versus the theoretical one, for the PT potential in the false vacuum at $\varphi_{0}=-2$. Here the best fit has $m=0.72$ and $C_{0}=-0.016$.
In order to determine which it is, note that we have already seen that $C_{0}$ is independent of the particular model, i.e. it is the same in the interacting PT theory as in the free-field theory with the SHO potential of Eq. (68). Moreover, when we increase the size of the Hilbert space $C_{0}$ begins to asymptote to zero. For example we find $C_{0}\approx 0.008$ when the number of quadrature points is increased from 20 to 30 (although it is computationally too demanding to go to a much more finely discretized quadrature). Therefore we can reasonably conclude that $C_{0}$ is also an effect of the truncated and discretized Hilbert space, and that both $m$ and $C_{0}$ will approach the idealised QFT values in the limit of infinite dimensionality of the qumode quadrature.
IV False vacuum decay in (1+1)-dimensional QFT
IV.1 The objective: tunnelling solutions
@@
$\partial^{2}_{r}\varphi+r^{-1}\partial_{r}\varphi\penalty 10000\ =\penalty 10000\ \partial_{\phi}{\mathscr{V}}\penalty 10000\ .$ (69)
This now resembles the classical equation of motion of a ball rolling in the inverted potential $\mathscr{V}$, with $r$ playing the role of time and with the $r^{-1}\partial_{r}\varphi$ term providing a friction that decays with time. Thus, the solution with the correct boundary conditions can be found by starting the ball close to the true vacuum at $\varphi\approx\varphi_{0}$ at ‘time’ $r=0$ (corresponding to the centre of the instanton) and letting it come to a halt at $r\to\infty$ precisely at the false vacuum at $\varphi=-\varphi_{0}$. This requires a fine-tuning of the starting value of $\varphi$, numerically solving the one dimensional differential equation in Eq. (69), and employing the overshoot-undershoot method described in Ref. [2] to satisfy the boundary conditions. The result for the quartic potential in Fig. 2 with $\lambda=0.5$, $\varphi_{0}=2$ and $\varepsilon=0.02$ is the bubble profile shown in Fig. 14. Although it is not possible to find a closed form for the radially symmetric solution, it can be very accurately modelled by the following function:
$\varphi(r)\penalty 10000\ \approx\penalty 10000\ -\varphi_{0}\,\tanh\beta(r-r_{0})$ (70)
@@
$\varphi(x,t)\penalty 10000\ \approx\penalty 10000\ -\,\varphi_{0}\,\tanh\beta(\sqrt{x^{2}-t^{2}}-r_{0})\penalty 10000\ .$ (71)
@@
$\varphi(x,t>|x|)\penalty 10000\ \approx-\varphi_{0}\,(1+2e^{-2\beta r_{0}}\cos{2\beta|t^{2}-x^{2}|})\penalty 10000\ .$ (72)
@@
$\sigma$ $\penalty 10000\ =\penalty 10000\ \int dr\frac{\dot{\varphi}^{2}}{2}+V$
$=\int_{-\varphi_{0}}^{\varphi_{\rm esc}}d\varphi\sqrt{2{\mathscr{V}}}\penalty 10000\ =\penalty 10000\ S_{1}\penalty 10000\ ,$ (73)
@@
$r_{c}\penalty 10000\ =\penalty 10000\ \frac{S_{1}}{2\varphi_{0}\varepsilon}\penalty 10000\ .$ (74)
@@
As mentioned in the introduction the rate of bubble nucleation per unit space-time volume is exponentially sensitive to the Euclidean action associated with the tunnelling solution. Specifically, in the thin-wall approximation it follows the Arrhenius-like behaviour $\Gamma\penalty 10000\ =\penalty 10000\ \kappa\,e^{-S_{1}}$, where $\kappa$ is a prefactor that accounts for quantum fluctuations around the instanton solution (which is notoriously difficult to evaluate). This exponential suppression means that even small increases in $S_{1}$ can lead to a dramatic decrease in the nucleation rate. Thus the value of $S_{1}$ effectively encodes how “difficult” it is for the system to tunnel from the false vacuum to the true vacuum, with lower $S_{1}$ corresponding to exponentially higher nucleation probabilities and faster phase transitions.
This then is the paradigm for false vacuum decay that was laid down in Refs. [2, 3]. Before considering the equivalent evolution on our qumode lattice, we should emphasize that this ‘classical $O(2)$ bounce’ picture provides a classical core for a process that has many subtleties.
To expand on this point (and also for later comparison), let us generate the entire nucleation and growth process for a single bubble (formed at time $t_{0}=20$), by numerically solving the PDEs. In Fig. 13 we show the numerical solution to the PDE for the $\tanh q^{\ell}$ potential of Eq. (24) with $\lambda=\mu=1$, $\ell=2$ and parameters $\varphi_{\rm fv}=-2$, $\varphi_{\rm tv}=0.7$, $\varepsilon=0.3$.
The solution is found by evolving from the bubble wall profile $\varphi(t_{0},x)$, determined by the aforementioned shooting method. This along with $\dot{\varphi}(t_{0},x)=0$ is our Cauchy data. We then evolve backwards for $t<t_{0}$ in Euclidean time, and forwards in Lorentzian time. Thus the hyperbolic evolution after the bubble has nucleated is effectively the analytic continuation of the Euclidean $O(2)$ bounce.
However, despite this nice global picture of the whole process, it is important to realise that the smooth Euclidean cap is not what one expects to see in pure Lorentzian time (i.e. in real time). Neither does the bubble simply ‘pop into existence’ all at once, but rather it first ‘appears’ where the negative eigenmode of the $O(2)$ bounce has support, which is at the walls. This real-time dynamics is the quantum mechanical aspect of the tunnelling process that is most hard to capture. Thus while the $O(2)$ bounce prescription gives a useful classical core for understanding the tunnelling process it tells us little about these more subtle aspects.
Figure 13: An ideal instanton tunnelling from the metastable QFT vacuum. The system is initialised in the metastable QFT vacuum with the $\tanh q^{\ell}$ potential of Eq. (24) with $\lambda=\mu=1$, $\ell=2$ and parameters $\varphi_{\rm fv}=-2$, $\varphi_{\rm tv}=0.7$, $\varepsilon=0.3$. The dashed white line marks the time where the bubble appears. Below this line time is imaginary: i.e. we analytically continue to Euclidean space-time which caps off the cone with a hemisphere.
Indeed there are many other subtle issues surrounding the dynamics of these processes which have been discussed subsequently, and which are not fully understood. For example, whether or not the bubble “runs away” to light speed or instead reaches a terminal velocity depends on the friction generated by its interactions with the medium it is travelling through, and is a complicated and model-dependent question (see [80, 81, 82, 83] and for a recent review see [84]). This of course is one of the motivations for studying the process numerically.
@@
IV.2 Results: tunnelling in (1+1)-dimensional scalar theory
To validate the model, we simulate the real-time dynamics of the system in three distinct scenarios. First, we consider “pre-formed” bubbles, initialising the system with a true-vacuum bubble with an initial radius $r_{0}$. Second, we study phase transitions in the full system by examining bubble nucleation arising from quantum fluctuations.
IV.2.1 Isolated bubbles: expansion and collapse
@@
Figure 15: In 15(a) we show a collapsing subcritical bubble in the Pöschl-Teller potential of Eq. (26) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\alpha=1$, $\gamma=3/2$ and $\varepsilon=0.25$. Fig. 15(b) shows an expanding bubble in the potential of Eq. (24) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\lambda=1$, and $\varepsilon=0.1$, which shows the bubble walls following a hyperbolic trajectory that reaches a terminal speed of 0.72. Note that we do not impose periodic boundary conditions, and the bubble walls “bounce” off the sides of the lattice elastically.
To begin, we examine the real-time dynamics of pre-formed bubbles with an initial radius, $r_{0}$. Here our initial state is taken to be sitting in the metastable “product of qumodes” vacuum, that is the quantum mechanical vacuum determined with no gradient-squared kinetic terms. As we have discussed, this is an approximation that is useful if long wavelength correlations are not important, which is likely to be the case when the bubble is already formed, in which case the evolution is expected to be more local and classical in nature.
@@
Fig. 15 presents the results from the qumode tensor network for bubbles of different initial radii. The first scenario, shown in Fig. 15(a), considers the collapse of a subcritical bubble in the Pöschl-Teller potential from Eq. (26) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\alpha=1$, $\gamma=3/2$ and $\varepsilon=0.25$. In this example, we observe that the bubble fails to reach the critical radius. There are two interesting aspects of the collapse. The first phenomenon that can be seen has to do with the way the collapse occurs. We can clearly see the emission of several wavepackets from the walls. Averaged over time these emissions generate friction on the walls. Here, however, we see in addition that such emission (together with momentum conservation) is a necessary component of bubble collapse. This is because, in one spatial dimension, a single bubble wall gains no local energetic advantage if it moves to collapse the bubble, whereas in higher dimensions there is a classical energy balance between surface tension and vacuum energy, which the surface tension wins if the bubble has a subcritical radius. On a local patch of wall we see this as a nett force accelerating the patch towards the centre of the bubble if the bubble is subcritical. This classical effect is absent in one dimension because the energetic cost of surface tension is independent of the bubble radius. The second interesting aspect of Fig. 15(a) is related. It is the fact that the final moments of collapse, in which the whole bubble “flops” back to the false vacuum, would correspond to super-luminal wall velocity. This in turn implies that the entanglement on the qumode lattice must spread over distances of at least the radius of the collapsing bubble at that instant in time, in order for the collapse to happen.
@@
Figure 15(b) shows the expansion of a bubble in the potential from Eq. (24) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\lambda=1$ and $\varepsilon=0.1$. In contrast with Fig. 15(a), this bubble is supercritical and we observe the phenomena outlined in Sec. IV.1. As the bubble’s radius is greater than the critical bubble radius, the walls expand. The bubble expansion is entirely driven by internal pressure caused by the energy gain from increasing the volume of the true vacuum. This acceleration is relatively local, and long-range entanglement plays very little role in the behaviour. As a result, the walls accelerate along hyperbolic trajectories which asymptote to a constant terminal velocity. However the terminal velocity is subluminal ($v_{\rm term}\approx 0.72$), due to the friction mentioned above.
@@
Note that we do not impose periodic boundary conditions and this results in the bubble wall colliding elastically with the sides of the lattice. We can in principle use this phenomenon to measure the mass of the bubble walls. We observe that after the bounce the walls are projected a distance of $d_{\rm bounce}\approx 20$ into the bulk of the lattice before being repelled back by the pressure, $\varepsilon$. (This relatively short distance also tells us that friction has indeed resulted in a terminal subluminal speed.) We may make a rough estimate for the mass of the wall by neglecting the effect of friction after the bounce and taking a non-relativistic approximation. Equating the initial kinetic energy with the maximum potential energy due to the pressure gives
$m_{\rm wall}\penalty 10000\ \approx\penalty 10000\ \frac{2\varepsilon d_{\rm bounce}}{v_{\rm term}^{2}}\penalty 10000\ \approx\penalty 10000\ 7.7\penalty 10000\ .$ (75)
Performing the integral in Eq. (IV.1) numerically we find that the actual wall mass should be $m_{\rm wall}\penalty 10000\ =\penalty 10000\ S_{1}\penalty 10000\ \approx\penalty 10000\ 2.74$ for this set of parameters, so this estimate is in the right ball-park. It could potentially be improved by taking into account friction.
IV.2.2 Phase transitions